OpenStudy (anonymous):

Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 i got this wrong

OpenStudy (anonymous):

Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 and this @YanaSidlinskiy

OpenStudy (anonymous):

My teacher didnt explain why though and i didnt fully understand. @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Did I help you with these before??

OpenStudy (anonymous):

yes different problem but same material.

OpenStudy (anonymous):

y = - 2x + 2 that was my answer.

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@mathmale